OpenStudy (calculusxy):

How to find the value of x in x^4 = 4 using a scientific calculator?

OpenStudy (calculusxy):

@phi @Directrix

OpenStudy (phi):

depends. there are 4 roots. but in your head you can take the square root of both sides to get \[ x^2= \pm 2 \] and if you take the positive 2, you get \( x= \pm \sqrt{2} \) which is 2 of the 4 roots.

OpenStudy (calculusxy):

i actually need to find the decimal form of it.

OpenStudy (calculusxy):

but how can i use guess and check effectively (w/o it being too time-consuming) to solve for x? @phi

OpenStudy (jdoe0001):

use the rational exponent button from the calculator \(\Large x^4=4\implies x=\sqrt[4]{4^1}\implies x=4^{\frac{1}{4}}\)

OpenStudy (calculusxy):

i am asking this question because it is related to (i think) another question. here it is: there are originally 1000 bacteria in a culture, and 4 hours later there are 4000. find the rate of increase per hour of the bacteria.

OpenStudy (calculusxy):

exponential growth: y = \(ab^x\) 4000 = 1000\(x^4\) 4000/1000 = (1000\(x^4\))/1000 4 = \(x^4\)

OpenStudy (calculusxy):

i am just learning algebra 1 and my teacher prefers for me not to use the other sophisticated formula of \(A = Pe^{r(t)}\)

OpenStudy (jdoe0001):

ok

OpenStudy (calculusxy):

so how can i solve that problem?

OpenStudy (calculusxy):

@jdoe0001

OpenStudy (phi):

your calculator can do roots do 4^(¼)